OpenStudy (anonymous):

What is the eccentricity of the ellipse given by the equation x^2/144+ y^2/64 = 1?

OpenStudy (anonymous):

Excentricity = c/a c = √(144 - 64) = √80 ≈ 9 a² = 144 ==> a = 12 c/a ≈ 9/12 ≈ 0.75

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

welcome